OpenStudy (anonymous):

please help me find particular integral of d2u/dx2 + u = cos (x-xo)

OpenStudy (solomonzelman):

what I see so far is: \(\large\color{black}{ \displaystyle \frac{ {\rm d}^2u}{ {\rm d}x^2}+u=\cos\left(x-x_0\right) }\)

OpenStudy (anonymous):

yes it is my que

OpenStudy (anonymous):

answer please

OpenStudy (anonymous):

how u find up what method is used...the ans that i have is (x-xo)/2 sin(x-xo)

OpenStudy (unklerhaukus):

Have you found the homogenous solution?

OpenStudy (anonymous):

no,i want to find in homogeneous solution

OpenStudy (unklerhaukus):

The homogenous solution solves \[u''+u=0\]

OpenStudy (anonymous):

ok if i ask u to solve up for d2u/dx2+u =exp(i(x-xo)) where i+iota

OpenStudy (unklerhaukus):

Variation of parameters?

OpenStudy (anonymous):

means?

OpenStudy (unklerhaukus):

Variation of Parameters, is one method to find the particular integral. But if you haven't learnt it, there are other methods.

OpenStudy (anonymous):

ok solve it by any method....but ans should be (x-xo)/2 cos(x-xo)+i sin(x-xo)